Solve each equation. Check the solutions.
2x-3=x-3

Answer by Edwin McCravy(20056)


 2x - 3 = x - 3

Subtract x from both sides:

 2x - 3 = x - 3
 -x      -x
----------------
  x - 3 =    -3

Add 3 to both sides

  x - 3 = -3
     +3   +3
 ------------
  x     =  0

To check the solution, substitute 0 for x
in the original equation:

   2x - 3 = x - 3
 2(0) - 3 = (0) - 3
    0 - 3 = 0 - 3
       -3 = -3

So it's right! 

Edwin